Stat215 Quiz 1 Spring 2019 For each variable below determine the variable type (continuous numeric, discrete numeric, categorical ranked, cat egorical unranked). 1. Height of a randomly selected building in Morgantown 2. The number of books a randomly selected adult owns 3. The color of the backpack of a randomly selected student 4. The cell phone number of a randomly selected US citizen 5. The weight of a randomly Wvu student Use the data that follows to answer questions 6-8 x: 19, 34, 28, 52, 47 6. mean 7. median 8. standard deviation
Answer #1

Answer :- [ 1 ] = continuous numeric

Answer :- [ 2 ] = discrete numeric

Answer :- [ 3 ] = categorical ranked

Answer :- [ 4 ] = discrete numeric

Answer :- [ 5 ] = continuous numeric

Data given :- 19, 34, 28,52,47

6) Mean = sum/n = 180/5 = 36

7) Median = 34 (middle most value )

8) Sample Standard deviation = sqrt ( ( sum(Xi - mean )² ) / n-1 ) =

SD = 734 / 4 = 12.12